# Semiclassical limit of resonance states in chaotic scattering

### v1

@@ Abstract
Resonance states in quantum chaotic scattering systems have a multifractal structure that depends on their decay rate. We show how classical dynamics describes resonance states of all decay rates in the semiclassical limit. This result for chaotic scattering systems corresponds to the well-established quantum ergodicity for closed chaotic systems. Specifically, we generalize Ulam’s matrix approximation of the Perron-Frobenius operator, giving rise to conditionally invariant measures of various decay rates. There are many matrix approximations leading to the same decay rate and we conjecture a criterion for selecting the one relevant for resonance states. Numerically, we demonstrate that resonance states in the semiclassical limit converge to the selected measure. Example systems are a dielectric cavity, the three-disk scattering system, and open quantum maps.
Introduction.—The structure of eigenstates in closed quantum systems, which in the classical limit are ergodic, is described by the semiclassical eigenfunction hypothesis [1, 2, 3] and the quantum ergodicity theorem [4, 5, 6, 7, 8, 9]. In the semiclassical limit almost all eigenstates are uniform on the energy surface in phase space.
In contrast, in quantum scattering systems with chaotic dynamics in the classical limit [10, 11, 12, 13], the structure of resonance states is much more complex [14, 15, 16], see Fig. 1. A recent factorization conjecture states that resonance states are composed of a universal factor given by a complex Gaussian random wave model and a factor of classical origin giving a multifractal structure depending on the decay rate [17, 18, 19]. Indeed, it was proven by Nonnenmacher and Rubin that in the semiclassical limit chaotic resonance states are described by some conditionally invariant measures [20]. Such measures are invariant under classical dynamics up to a change of their norm due to escape [21].
@@
Figure 1: (a) Visual comparison between the average over 500 resonance states of a dielectric cavity of limaçon shape near $\mathrm{Re}\,kR_{\text{cav}}=2000$ with decay rate near $\gamma=0.053$ (top) and the proposed conditionally invariant measure based on ray dynamics (bottom). (b) Comparison on boundary phase space.
@@
In this paper we construct the conditionally invariant measure that describes the structure of chaotic resonance states of a given decay rate in the semiclassical limit. To this end we generalize Ulam’s matrix approximation of the Perron-Frobenius operator. There are many matrix approximations and we conjecture a criterion for selecting the one relevant for resonance states, giving the desired conditionally invariant measure. Numerically, we demonstrate that resonance states in the semiclassical limit converge to the selected measure. A visual comparison is shown for a dielectric cavity (partial escape) in Fig. 1. Further example systems are the three-disk scattering system (full escape), and quantum maps with partial and full escape.
Ulam’s method.—The time evolution of densities in the phase space of a dynamical system is governed by the Perron-Frobenius operator [22]. A matrix approximation of the operator goes back to Ulam [54] and has found many applications [55, 56, 57, 58, 59, 60, 61]. In the context of scattering systems it was applied to maps with full escape [62] and dielectric cavities [31]. There it generates the natural conditionally invariant measure with the natural decay rate. We first describe this approach, often called Ulam’s method, which is generalized to arbitrary decay rates below.
One partitions phase space into $n$ disjoint cells $\{A_{1},\dots,A_{n}\}$, typically a grid of equally sized boxes. For simplicity, we first consider a time-discrete map $T$ on phase space. It defines for each cell $A_{i}$ subregions $A_{ji}\subset A_{i}$, which are mapped to cells $A_{j}$ under $T$. In other words, each subregion $A_{ji}$ is defined by,
@@
where $\mu_{\mathcal{L}}$ is the uniform Lebesgue measure on phase space. We denote it as the Lebesgue transition matrix $P^{\mathcal{L}}$ to distinguish it from more general transition matrices $P$ to be introduced below. Each element is the ratio of the Lebesgue measure of the subregion $A_{ji}$ to the Lebesgue measure of the whole cell $A_{i}$. From the above definitions follows $\sum_{j}P^{\mathcal{L}}_{ji}=1$, as required for a transition matrix. The matrix $P^{\mathcal{L}}$ is Ulam’s matrix approximation of the Perron-Frobenius operator. Numerically, it is determined by the fraction of trajectories uniformly started within cell $A_{i}$ that goes to cell $A_{j}$. We mention that for closed systems the right leading eigenvector, $\sum_{i}P^{\mathcal{L}}_{ji}\,\mu_{i}=\mu_{j}$, gives a coarse-grained invariant measure $\mu_{i}=\mu(A_{i})$ for all cells $A_{i}$ of the partition [54, 22]. Note that we use a notation for the order of indices common in physics, as e.g. in Refs. [62, 31].
In a scattering system with partial escape one approximates the reflectivity by a matrix $R$. Each element $R_{ji}$ is representative for the reflectivity of the transition $A_{i}\rightarrow A_{j}$. In a system with full escape some of the $R_{ji}$ will be zero. Combining a transition matrix $P$ with the reflectivity matrix $R$ by element-wise multiplication leads to a matrix $P_{ji}R_{ji}$ with the eigensystem,
@@
Here the right leading eigenvector defines a coarse-grained conditionally invariant measure $\mu_{i}=\mu(A_{i})$ for all cells $A_{i}$ of the partition. The leading eigenvalue $\text{e}^{-\gamma}$ gives the decay rate $\gamma$ of this measure. Note that $\mu_{i}\geq 0$ for all $i$ is ensured by the Perron-Frobenius theorem [63] for the leading eigenvector. Subleading solutions have no relevance here. This approach was introduced in Ref. [62] for maps with full escape using the Lebesgue transition matrix $P^{\mathcal{L}}$.
For time-continuous systems one also has to consider the transition time [16], e.g. between reflections with a billiard boundary, while in Eq. (3) the transition time was implicitly set to 1. The transition time is approximated by a matrix $t$, where each element $t_{ji}$ is representative for the transition $A_{i}\rightarrow A_{j}$. In this case, the exponential factor in Eq. (3) has to be placed on the left hand side, yielding the eigensystem,
@@
with leading eigenvalue $1$. Here the unknown decay rate $\gamma$ has to be adjusted, such that the leading eigenvalue of the matrix $P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}$ (multiplied element-wise) is indeed $1$. This can be done iteratively [64]. This approach (without adjusting $\gamma$) was introduced in Ref. [31] for dielectric cavities using the Lebesgue transition matrix $P^{\mathcal{L}}$.
In the limit of increasingly finer partitions of phase space the number of cells increases, $n\rightarrow\infty$. One desires that in this limit the coarse-grained conditionally invariant measure converges, which is mathematically an open question. Numerically, using the Lebesgue transition matrix $P^{\mathcal{L}}$ in Eqs. (3) or (4) one finds convergence to the natural measure $\mu_{\text{nat}}$ and the natural decay rate $\gamma_{\text{nat}}$, respectively [62, 31]. Note that the natural measure $\mu_{\text{nat}}$ is typically not determined using a matrix approximation. Instead a long-term time evolution of any smooth density with the Perron-Frobenius operator is done, which is numerically implemented by iterating trajectories and their intensities [14, 23, 16].
Transition matrices for arbitrary decay rates.—The Lebesgue transition matrix $P^{\mathcal{L}}$, Eq. (2), is based on the restriction that the measure of the subregions $A_{ji}$ of a cell $A_{i}$ is given by the uniform Lebesgue measure. This assumption limits the possible coarse-grained measures to the natural measure decaying with $\gamma_{\text{nat}}$.
@@
Selection criterion for transition matrix.—Which of the above infinitely many coarse-grained conditionally invariant measures for a given decay rate $\gamma$ is relevant for describing quantum mechanical resonance states with this decay rate? We conjecture that it is the measure emerging from the transition matrix $P$ that is closest to $P^{\mathcal{L}}$, in the sense of minimizing the Kullback-Leibler divergence from $P$ to $P^{\mathcal{L}}$,
@@
Figure 2: Cell $A_{i}$ of the partition and its subregions $A_{ji}\subset A_{i}$, Eq. (1). (a) Uniform density giving Lebesgue transition matrix $P^{\mathcal{L}}$. (b) Non-uniform density giving a general transition matrix $P$.
For a better intuition it helps to relate the elements $P_{ji}$ of a transition matrix to a density distribution within a cell $A_{i}$, see Fig. 2. For the Lebesgue transition matrix $P^{\mathcal{L}}$, Eq. (2), the measure $\mu_{\mathcal{L}}(A_{ji})$ of each subregion $A_{ji}$ is obtained by integrating a uniform density. Instead, for a general transition matrix $P$, Eq. (5), we allow for a non-uniform density within the cell. Such $P$ give rise to conditionally invariant measures with other decay rates. The selection criterion chooses among the non-uniform densities, which gives rise to the desired decay rate, the one closest to a uniform density.
@@
Example systems.—We will use four example systems to demonstrate for the selected measure (i) the convergence in the limit $n\rightarrow\infty$ of finer partitions and (ii) the agreement with resonance states in the semiclassical limit. These examples cover the cases of partial and full escape for a 2D billiard and a map in each case:
@@
For details on these systems see [64]. We provide Python code to compute $P^{\mathcal{L}}$, $P$, and $\mu$ [64]. In all cases the elements of $P^{\mathcal{L}}$ are determined using $10^{4}$ trajectories per cell of the partition (started on a uniform grid) for up to $n=3200^{2}\approx 10^{7}$ cells.
(i) Convergence in the limit $n\rightarrow\infty$.—We numerically demonstrate the convergence of the coarse-grained measure in the limit $n\rightarrow\infty$ of a finer partition. To this end we compare the coarse-grained measures for $n$ and for $n/4$ cells by integrating them on a $50\times 50$ grid and using the Jensen-Shannon divergence [66] (its square root being a metric). Figure 3 shows that with increasing $n$ the Jensen-Shannon divergence converges to zero. Furthermore, the ratio of two consecutive Jensen-Shannon divergences is bounded from above by a value smaller than $1$. If this continues to hold for increasing $n$, the coarse-grained measures are a contractive sequence, therefore a Cauchy sequence and thus converge in the limit $n\rightarrow\infty$. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma\geq\gamma_{\text{nat}}$ (full escape) for the billiard systems and maps.
@@
(ii) Comparison with resonance states in the semiclassical limit.—We compare Husimi representations of quantum resonance states near some decay rate $\gamma$ to the corresponding proposed measure from the finest partition. In order to be most sensitive, we use the average over 500 resonance states of similar decay rate. The visual comparison is demonstrated in Fig. 1 for the dielectric cavity at one decay rate, for further examples see [64]. Quantitatively, Fig. 4 shows that going further towards the semiclassical limit the Jensen-Shannon divergence (evaluated on a $50\times 50$ grid) converges to zero. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma\geq\gamma_{\text{nat}}$ (full escape) for the billiard systems and maps. This convergence to zero is in contrast to previous approximate approaches for conditionally invariant measures [53, 36], which describe resonance states quite well, but not perfectly [36, 19]. We attribute different absolute values and convergence speeds to properties of the multifractal resonance states. These numerical results for various types of scattering systems give strong support for the conjectured selection criterion for the transition matrix and the corresponding measure.
Remarks.—Let us stress, that for individual resonance states we find just as well a Jensen-Shannon divergence converging to zero [64]. However, one observes a much larger Jensen-Shannon divergence due to their fluctuations. This makes individual resonance states a much less sensitive test for the quality of a conditionally invariant measure [36].
The analysis is presented for right resonance states. It can be straightforwardly extended to left resonance states of billiards or maps with escape.
@@
Outlook.—It is desirable to find a semiclassical derivation for our selection criterion of the transition matrix leading to the proposed measures, as it is possible for locally randomized systems [64].
@@
Furthermore, it would be interesting to extend the set of examples to systems with scattering in smooth potential, e.g. models of the chaotic ionization of atoms and of resonances in chemical reactions [68, 69]. Also the relation to the recently found properties of Schur eigenstates [52] needs to be investigated, as well as the relation to quantum and classical channels in bipartite many-body systems [70].
@@ Acknowledgements.
We are grateful for discussions with E. Altmann, A. Bäcker, S. Barkhofen, K. Clauß, T. Harayama, M. Richter, P. Schlagheck, A. Shudo, G. Tanner, T. Tél, and T. Weich. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 262765445. The authors gratefully acknowledge the computing time on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR.

### v2

@@ Abstract
Resonance states in quantum chaotic scattering systems have a multifractal structure that depends on their decay rate. We show how classical dynamics describes this structure for all decay rates in the semiclassical limit. This result for chaotic scattering systems corresponds to the well-established quantum ergodicity for closed chaotic systems. Specifically, we generalize Ulam’s matrix approximation of the Perron-Frobenius operator, giving rise to conditionally invariant measures of various decay rates. There are many matrix approximations leading to the same decay rate and we conjecture a criterion for selecting the one relevant for resonance states. Numerically, we demonstrate that resonance states in the semiclassical limit converge to the selected measure. Example systems are a dielectric cavity, the three-disk scattering system, and open quantum maps.
Introduction—The structure of eigenstates in closed quantum systems, which in the classical limit are ergodic, is described by the semiclassical eigenfunction hypothesis [1, 2, 3] and the quantum ergodicity theorem [4, 5, 6, 7, 8, 9]. In the semiclassical limit almost all eigenstates are uniform on the energy surface in phase space.
In contrast, in quantum scattering systems with chaotic dynamics in the classical limit [10, 11, 12, 13], the structure of resonance states is much more complex [14, 15, 16], see Fig. 1. A recent factorization conjecture states that resonance states are composed of a universal factor given by a complex Gaussian random wave model and a factor of classical origin giving a multifractal structure depending on the decay rate [17, 18, 19]. Indeed, it was proven by Nonnenmacher and Rubin that in the semiclassical limit chaotic resonance states are described by some conditionally invariant measures [20]. Such measures are invariant under classical dynamics up to a change of their norm due to escape [21, 22].
@@
Figure 1: (a) Visual comparison between the average over 500 resonance states of a dielectric cavity of limaçon shape near $\mathrm{Re}\,kR_{\text{cav}}=5000$ with decay rate near $\gamma=0.053$ (top) and the proposed conditionally invariant measure based on ray dynamics (bottom). (b) Comparison on boundary phase space.
@@
In this Letter we construct the conditionally invariant measure that describes the structure of chaotic resonance states of a given decay rate in the semiclassical limit. To this end we generalize Ulam’s matrix approximation of the Perron-Frobenius operator. There are many matrix approximations and we conjecture a criterion for selecting the one relevant for resonance states, giving the desired conditionally invariant measure. Numerically, we demonstrate that resonance states in the semiclassical limit converge to the selected measure. A visual comparison is shown for a dielectric cavity (partial escape) in Fig. 1. Further example systems are the three-disk scattering system (full escape), and quantum maps with partial and full escape.
Ulam’s method—The time evolution of densities in the phase space of a dynamical system is governed by the Perron-Frobenius operator [23]. A matrix approximation of the operator goes back to Ulam [55] and has found many applications [56, 57, 58, 59, 60, 61, 62]. In the context of scattering systems it was applied to maps with full escape [63] and dielectric cavities [32]. There it generates the natural conditionally invariant measure with the natural decay rate. We first describe this approach, often called Ulam’s method, which is later generalized to arbitrary decay rates.
One partitions phase space into $n$ disjoint cells $\{A_{1},\dots,A_{n}\}$, typically a grid of equally sized boxes. For simplicity, we first consider a time-discrete map $T$ on phase space. It defines for each cell $A_{i}$ subregions $A_{ji}\subset A_{i}$, which are mapped to cells $A_{j}$ under $T$. In other words, each subregion $A_{ji}$ is defined by
@@
where $\mu_{\mathcal{L}}$ is the uniform Lebesgue measure on phase space. We denote it as the Lebesgue transition matrix $P^{\mathcal{L}}$ to distinguish it from more general transition matrices $P$ to be introduced below. Each element is the ratio of the Lebesgue measure of the subregion $A_{ji}$ to the Lebesgue measure of the whole cell $A_{i}$. From the above definitions follows $\sum_{j}P^{\mathcal{L}}_{ji}=1$, as required for a transition matrix. The matrix $P^{\mathcal{L}}$ is Ulam’s matrix approximation of the Perron-Frobenius operator. Numerically, it is determined by the fraction of trajectories uniformly started within cell $A_{i}$ that goes to cell $A_{j}$. We mention that for closed systems the right leading eigenvector, $\sum_{i}P^{\mathcal{L}}_{ji}\,\mu_{i}=\mu_{j}$, gives a coarse-grained invariant measure $\mu_{i}=\mu(A_{i})$ for all cells $A_{i}$ of the partition [55, 23]. Note that we use a notation for the order of indices common in physics, as, e.g., in Refs. [63, 32].
In a scattering system with partial escape one approximates the reflectivity by a matrix $R$. Each element $R_{ji}$ is representative for the reflectivity of the transition $A_{i}\rightarrow A_{j}$. In a system with full escape some of the $R_{ji}$ will be zero. Combining a transition matrix $P$ with the reflectivity matrix $R$ by elementwise multiplication leads to a matrix $P_{ji}R_{ji}$ with the eigensystem,
@@
Here the right leading eigenvector defines a coarse-grained conditionally invariant measure $\mu_{i}=\mu(A_{i})$ for all cells $A_{i}$ of the partition. The leading eigenvalue $\text{e}^{-\gamma}$ gives the decay rate $\gamma$ of this measure. Note that $\mu_{i}\geq 0$ for all $i$ is ensured by the Perron-Frobenius theorem [64] for the leading eigenvector. Subleading solutions have positive and negative entries. Thus they cannot be interpreted as measures and are of no relevance here. The above approach was introduced in Ref. [63] for maps with full escape using the Lebesgue transition matrix $P^{\mathcal{L}}$.
For time-continuous systems one also has to consider the transition time [16], e.g. between reflections with a billiard boundary, while in Eq. (3) the transition time was implicitly set to 1. The transition time is approximated by a matrix $t$, where each element $t_{ji}$ is representative for the transition $A_{i}\rightarrow A_{j}$. In this case, the exponential factor in Eq. (3) has to be placed on the left-hand side, yielding the eigensystem,
@@
with leading eigenvalue $1$. Here the unknown decay rate $\gamma$ has to be adjusted, such that the leading eigenvalue of the matrix $P_{ji}R_{ji}\text{e}^{\gamma t_{ji}}$ (multiplied elementwise) is indeed $1$. This can be done iteratively [65]. This approach (without adjusting $\gamma$) was introduced in Ref. [32] for dielectric cavities using the Lebesgue transition matrix $P^{\mathcal{L}}$.
In the limit of increasingly finer partitions of phase space the number of cells increases, $n\rightarrow\infty$. One desires that in this limit the coarse-grained conditionally invariant measure converges, which is mathematically an open question. In general, the chosen partition and functional space influence the convergence, see the recent Ref. [62] and references therein. Numerically, using the Lebesgue transition matrix $P^{\mathcal{L}}$ in Eqs. (3) or (4) one finds convergence to the natural measure $\mu_{\text{nat}}$ and the natural decay rate $\gamma_{\text{nat}}$, respectively [63, 32]. Note that the natural measure $\mu_{\text{nat}}$ is typically not determined using a matrix approximation. Instead a long-term time evolution of any smooth density with the Perron-Frobenius operator is done, which is numerically implemented by iterating trajectories and their intensities [14, 24, 16].
Transition matrices for arbitrary decay rates—The Lebesgue transition matrix $P^{\mathcal{L}}$, Eq. (2), is based on the restriction that the measure of the subregions $A_{ji}$ of a cell $A_{i}$ is given by the uniform Lebesgue measure. This assumption limits the possible coarse-grained measures to the natural measure decaying with $\gamma_{\text{nat}}$.
@@
Selection criterion for transition matrix—Which of the above infinitely many coarse-grained conditionally invariant measures for a given decay rate $\gamma$ is relevant for describing quantum mechanical resonance states with this decay rate? We conjecture that it is the measure emerging from the transition matrix $P$ that is closest to $P^{\mathcal{L}}$, in the sense of minimizing the Kullback-Leibler divergence from $P$ to $P^{\mathcal{L}}$,
@@
Figure 2: Cell $A_{i}$ of the partition and its subregions $A_{ji}\subset A_{i}$, Eq. (1). (a) Uniform density giving Lebesgue transition matrix $P^{\mathcal{L}}$. (b) Nonuniform density giving a general transition matrix $P$.
@@
For a better intuition it helps to relate the elements $P_{ji}$ of a transition matrix to a density distribution within a cell $A_{i}$, see Fig. 2. For the Lebesgue transition matrix $P^{\mathcal{L}}$, Eq. (2), the measure $\mu_{\mathcal{L}}(A_{ji})$ of each subregion $A_{ji}$ is obtained by integrating a uniform density. Instead, for a general transition matrix $P$, Eq. (5), we allow for a nonuniform density within the cell. Such $P$ give rise to conditionally invariant measures with other decay rates. The selection criterion chooses among the nonuniform densities giving the desired decay rate, the one closest to a uniform density.
@@
Example systems—We will use four example systems to demonstrate for the selected measure (i) the convergence in the limit $n\rightarrow\infty$ of finer partitions and (ii) the agreement with resonance states in the semiclassical limit. These examples cover the cases of partial and full escape for a 2D billiard and a map in each case:
@@
For details on these systems see [65]. We provide Python code to compute $P^{\mathcal{L}}$, $P$, and the proposed measure $\mu$ [65]. In all cases the elements of $P^{\mathcal{L}}$ are determined using $10^{4}$ trajectories per cell of the partition (started on a uniform grid) for up to $n=3200^{2}\approx 10^{7}$ cells.
(i) Convergence in the limit $n\rightarrow\infty$—We numerically demonstrate the convergence of the coarse-grained measure in the limit $n\rightarrow\infty$ of a finer partition. To this end we compare the coarse-grained measures for $n$ and for $n/4$ cells by integrating them on a $50\times 50$ grid and using the Jensen-Shannon divergence [67] (its square root being a metric). Figure 3 shows that with increasing $n$ the Jensen-Shannon divergence converges to zero. Furthermore, the ratio of two consecutive Jensen-Shannon divergences is bounded from above by a value smaller than $1$. If this continues to hold for increasing $n$, the coarse-grained measures are a contractive sequence, therefore a Cauchy sequence and thus converge in the limit $n\rightarrow\infty$. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma\geq\gamma_{\text{nat}}$ (full escape) for the billiard systems and maps.
@@
(ii) Comparison with resonance states in the semiclassical limit—We compare Husimi representations of quantum resonance states near some decay rate $\gamma$ to the corresponding proposed measure from the finest partition. In order to be most sensitive, we use the average over 500 resonance states of similar decay rate. The visual comparison is demonstrated in Fig. 1 for the dielectric cavity at one decay rate. Note that resonance states show multifractality on scales larger than a Planck cell only and we have to smooth the measure on this scale for the visual comparison in Fig. 1(b). For further examples see Supplemental Material [65].
Quantitatively, Fig. 4 shows that going further towards the semiclassical limit the Jensen-Shannon divergence (evaluated on a $50\times 50$ grid) converges to zero. This is demonstrated for various decay rates $\gamma\in[\gamma_{\text{nat}},\gamma_{\text{inv}}]$ (partial escape) and $\gamma\geq\gamma_{\text{nat}}$ (full escape) for the billiard systems and maps. This convergence to zero is in contrast to previous approximate approaches for conditionally invariant measures [54, 37], which describe resonance states quite well, but not perfectly [37, 19]. We attribute different absolute values and convergence speeds to properties of the multifractal resonance states. These numerical results for various types of scattering systems give strong support for the conjectured selection criterion for the transition matrix and the corresponding measure.
Remarks—Let us stress that for individual resonance states we find just as well a convergence to the proposed measure [65], as expected by the factorization conjecture [17, 18, 19]. However, one observes a much larger Jensen-Shannon divergence due to their fluctuations. This makes individual resonance states a much less sensitive test for the quality of a conditionally invariant measure [37].
The analysis is presented for right resonance states. It can be straightforwardly extended to left resonance states of billiards or maps with escape. This should allow for predicting the structure of the left-right Husimi representation [44].
@@
Outlook—It is desirable to find a semiclassical derivation for our selection criterion of the transition matrix leading to the proposed measures, as it is possible for locally randomized systems [65]. Numerically, it might be of interest to find an alternative method to compute the proposed measures based on iterating trajectories for long times, as it is common for the natural measure.
@@
Furthermore, it would be interesting to extend the set of examples to systems with scattering in smooth potential, e.g., models of the chaotic ionization of atoms and of resonances in chemical reactions [69, 70]. Also the relation to the recently found properties of Schur eigenstates [53] needs to be investigated, as well as the relation to quantum and classical channels in bipartite many-body systems [71].
@@ Acknowledgements.
Acknowledgments—We are grateful for discussions with E. Altmann, A. Bäcker, S. Barkhofen, K. Clauß, T. Harayama, M. Richter, P. Schlagheck, A. Shudo, G. Tanner, T. Tél, and T. Weich. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)—262765445. The authors gratefully acknowledge the computing time on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR.